# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ 3.2 Estimators of Thresholds
We then design a Growing-Window (GW) approach to estimating the unknown and stochastic thresholds. Unlike the Monte Carlo (MC) estimator that aggregates all historical data (see Appendix C), the GW estimator dynamically focuses on more recent observations, offering better adaptability to non-stationary environments while still preserving asymptotic consistency.
Growing-Window estimator
We introduce a sequence of window size $\{W_{t}\}_{t=1}^{T}$, growing with episode $t$. The GW estimator at episode $t$ computes threshold estimates using data from the most recent $W_{t}$ episodes as follows:
[…]
where $N_{h}^{(W_{t}),t}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ counts visits to state-action pair $(s,a)$ in the last $W_{t}$ episodes before episode $t$. As $t$ increases, the window size $W_{t}$ grows. We prove its asymptotic consistency as follows.
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}(s,a)\right|\leq\zeta_{h}^{(W_{t}),t}(s,a),$
where $\zeta_{h}^{(W_{t}),t}(s,a)=\min\left\{H,\sqrt{\frac{4H^{2}\ln(mSAHT/\delta)}{\max\{1,N_{h}^{(W_{t}),t}(s,a)\}}}\right\}$.
When there is no ambiguity, we simplify the notation $\hat{\alpha}_{i,h}^{(W_{t}),t}$ to $\hat{\alpha}_{i,h}^{t}$ and $\zeta_{h}^{(W_{t}),t}$ to $\zeta_{h}^{t}$. The proof is given in Appendix C.
@@
Fixed-window approaches may repeatedly discard valuable historical data, which hurts long-run convergence. Full-history methods may respond too slowly to changes. A Growing-Window approach preserves enough past information to ensure eventual convergence but can adapt to evolving conditions more rapidly – we desire this balance between statistical consistency and responsiveness to shifts in the environment or thresholds.
@@
Pessimistic and optimistic estimators
We also consider two ‘extreme’ estimation strategies, namely a conservative (pessimistic) approach and an optimistic estimation, defined as below,
$\overline{\alpha}_{i,h}^{t}(s,a):=\hat{\alpha}_{i,h}^{t-1}(s,a)+\zeta_{h}^{t-1}(s,a),\quad\underline{\alpha}_{i,h}^{t}(s,a):=\hat{\alpha}_{i,h}^{t-1}(s,a)-\zeta_{h}^{t-1}(s,a),$ (3)
for every pair $(s,a)$, step $h\in[H]$ and constraint $i\in[m]$.
Remark 3.
With probability at least $1-\delta$, any possible estimator $\hat{\alpha}_{i}$ satisfies inequality: $\underline{\alpha}_{i}\leq\hat{\alpha}_{i}\leq\overline{\alpha}_{i}$, for any constraint $i\in[m]$. This suggests that pessimistic and optimistic estimators are upper and lower boundaries of all possible estimators, respectively. Therefore, by analysing these pessimistic and optimistic estimators, we can effectively study and bound the performance of the GW estimator.
@@ 3.3 Feasibility and Strong Duality
A policy $\pi$ is defined to be pessimistic if the expected cumulative constraint is no less than the pessimistic threshold; i.e., $V_{g}^{\pi}\geq\overline{\alpha}$. Let $\mathcal{F}_{\textup{pes}}$ denote its feasibility set,
@@
Similarly, we define a policy $\pi$ to be optimistic if the expected cumulative constraint is greater than or equal to the optimistic threshold, i.e., $V_{g}^{\pi}\geq\underline{\alpha}$, and its corresponding feasibility set as below,
@@
From the definitions, pessimistic and optimistic policies are boundaries of all possible policies: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}(\omega)\subseteq\mathcal{F}_{\textup{opt}}$ with probability at least $1-\delta$.
Pessimistic and optimistic policies can be obtained from the following optimisation problems, respectively,
$\max_{q\in\mathcal{F}_{\text{pes}}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\overline{\alpha}(\omega)),$ (Pessimistic)
$\max_{q\in\mathcal{F}_{\text{opt}}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\underline{\alpha}(\omega)),$ (Optimistic)
Their strong dualities are characterised by the following lemmas.
@@
If there exists $q^{\pi_{1}}\in\mathcal{F}_{\text{pes}}$ and $\xi\in\mathbb{R}_{>0}^{m}$ such that $G^{T}q^{\pi_{1}}\geq\overline{\alpha}_{i}+\xi_{i}$ for all constraint $i\in[m]$, then we have $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{pes}}}\mathcal{L}(q,\lambda,\omega)=\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda}\mathcal{L}(q,\lambda,\omega).$
@@
If there exists $q^{\pi_{2}}\in\mathcal{F}_{\text{opt}}$ and $\xi^{\prime}\in\mathbb{R}_{>0}^{m}$ such that $G^{T}q^{\pi_{2}}\geq\underline{\alpha}_{i}+\xi^{\prime}_{i}$ for all constraint $i\in[m]$, then we obtain $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{opt}}}\mathcal{L}(q,\lambda,\omega)=\max_{q\in\mathcal{F}_{\textup{opt}}}\min_{\lambda}\mathcal{L}(q,\lambda,\omega).$
Each lemma requires the existence of a policy that strictly satisfies the corresponding thresholds. We illustrate these feasible sets and their corresponding feasible solutions under pessimistic, intermediate (thresholds derived from random sampling), and optimistic thresholds in Figure 1.
We set the Slater gaps under pessimistic and optimistic thresholds as follows, respectively,
$\rho:=(r^{T}q^{\pi^{\star}}-r^{T}q^{\pi_{1}})/\min_{i\in[m]}\xi_{i},$
$\rho^{\prime}:=(r^{T}q^{\pi^{\star}}-r^{T}q^{\pi_{2}})/\min_{i\in[m]}\xi^{\prime}_{i}.$
Remark 5.
As discussed, with probability at least $1-\delta$, pessimistic thresholds are upper bounds for all possible thresholds. If there exists a policy that strictly satisfies pessimistic thresholds, then it also satisfies all possible thresholds, ensuring that the strong duality holds.
@@ 3.4 A Primal-Dual Approach for CMDPs with Stochastic Thresholds
In this section, we propose SPOT (Stochastic Pessimistic-Optimistic Thresholding), a novel algorithm specifically tailored for CMDPs with stochastic thresholds. Unlike the closest algorithm OptPrimalDual-CMDPefroni2020exploration , which relies on fixed and fully known thresholds, SPOT is designed to address the more challenging and realistic scenarios where thresholds are both unknown and dynamically evolving over time.
@@
0:  number of episodes $T$, $\bf\lambda_{1}=0$, stepsizes $\eta_{t}$ and $\eta_{\lambda}$, and initial policy $\pi_{h}^{1}(a|s)=1/A$ for all $s$, $a$, $h$.
@@
9:      Execute $\pi_{t}$ via Algorithm 1 and update counters and empirical model (i.e., $\hat{\alpha}^{t},\hat{r}^{t},\hat{g}^{t},\hat{p}^{t}$) via Equations (2) and (7)
@@
For a given dual variable, the algorithm first updates the primal policy based on the estimated $Q$-values and subsequently refines the dual value using the estimated value. In the first stage (Line 2), it updates the optimistic estimators $\overline{r}^{t}$, $\overline{g}^{t}$ and $\overline{p}^{t}$. Then, it computes the estimator $\alpha_{\diamond}^{t}$ as prescribed in Equation (3). Importantly, depending on the problem context, estimator $\alpha_{\diamond}^{t}$ can be either pessimistic estimator $\overline{\alpha}^{t}$ or optimistic estimator $\underline{\alpha}^{t}$, thereby accommodating different degrees of uncertainty in the thresholds. Thus, at each episode $t$, it yields a new CMDP $\mathcal{M}_{t}:=(\mathcal{S},\mathcal{A},\overline{r}^{t},\overline{g}^{t},\alpha_{\diamond}^{t},\overline{p}^{t})$, which the algorithm then optimises via
$\max_{q}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}^{\prime}(q,\lambda,\omega):=\overline{r}^{T}q^{\pi}+\lambda^{T}(\overline{G}^{T}q^{\pi}-\alpha_{\diamond}(\omega)).$
@@
$\lambda_{t+1}=\Pi_{\mathcal{C}}\left(\lambda_{t}-\frac{1}{\eta_{\lambda}}(\overline{G}_{t-1}q^{\pi_{t}}-\alpha_{\diamond}^{t})\right),$ (4)
where $\overline{G}_{t-1}q^{\pi_{t}}$ corresponds to the value estimate $\hat{V}_{\overline{g}}^{\pi_{t}}$ and the projection set is defined as $\mathcal{C}:=\{\lambda\in\mathbb{R}^{m}:0\leq\lambda\leq\rho\bf 1\}$ with $\rho$ being the Slater gap defined in Section 3.3 and the projection operator $\Pi_{\mathcal{C}}:\mathbb{R}^{m}\to\mathcal{C}$ is defined by $\Pi_{\mathcal{C}}(x):=\arg\min_{y\in\mathcal{C}}\|y-x\|_{2}$. The feasibility proof for Equation (4) is presented in Appendix F. Finally, the policy $\pi_{t}$ is executed to collect new data, which is then used to update the threshold estimates and empirical models for subsequent iterations (Line 9).
@@
  • Stochastic and unknown thresholds vs. fixed thresholds. SPOT can handle thresholds that are unknown and random, going beyond the scope of fixed-threshold methods, which fit in more realistic scenarios.
  • Dual-mode thresholding for exploration and safety. SPOT introduces a novel dual-mode thresholding mechanism, enabling the selection of either pessimistic or optimistic thresholds based on the application’s risk tolerance: pessimistic thresholds ensure robust feasibility under worst-case assumptions, while optimistic thresholds enable more aggressive exploration when additional risk is acceptable. This twofold approach empowers SPOT to adapt to varying levels of constraint uncertainty.
  • Strong Performance Guarantees. By continuously refining its threshold estimates, SPOT can maintain performance guarantees that are the same order as fixed-threshold approaches, while maintaining adaptability to dynamic changes.
@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Let $S$, $A$, $m$, $H$, $T$ be the number of states, actions, constraints, horizon length, and episodes, respectively. Define $\rho$ as the Slater gap under pessimistic thresholds, $\gamma$ as the Growing-Window parameter, and $\mathcal{N}$ as the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability at least $1-\delta$, Algorithm 2 under pessimistic thresholds achieves the following regret bounds:
$\mathcal{R}_{T}(r)$ $\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+\sqrt{(1+m\rho)^{2}SAH^{4}T/\gamma}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right),$
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
Efroni et al. efroni2020exploration first introduce online methods that achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret and constraint violation in CMDPs; later, Ghosh et al. ghosh2024towards prove that $\tilde{\mathcal{O}}(\sqrt{T})$ regret and hard violation can be achieved by a primal-dual approach in tabular CMDPs. This paper eliminates the reliance on fixed thresholds, and achieves comparable guarantees.
@@
We present a proof sketch here, with the detailed proof provided in Appendix H.
@@
Decompose the Lagrangian $\mathcal{L}(\lambda)$, which equals the sum of the reward regret and the constraint violation scaled by dual variable as follows,
$\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda}_{(b)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(c)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(d)}+\underbrace{\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}}_{(e)}.$ (5)
where the notation appearing in the formula is defined in Appendix G. The decomposition shows that term (a) quantifies the cumulative error between the estimated policy and the optimal policy with $\lambda_{t}$, term (b) aggregates the errors in estimating the constraints and their thresholds, term (c) measures the gap between actual reward and the optimistic estimates and terms (d)-(e) capture the error from updating the dual variables.
@@
Upper bound each term in the decomposition. Bounding term (a) is one of major technical contributions in this paper. We obtain the following lemma, with the proof deferred to Appendix G.
Lemma 3 (Policy optimality of pessimistic policies).
Conditioned on the success event (formally defined in Appendix D), for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$. Under pessimistic thresholds, it holds
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left(\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}\right).$
@@
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to step $h$ at episode $t$. Suppose that for every $(t,h)$, it satisfies $\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h-1}^{t}]\right|\leq 2H.$ Then, for any for any confidence parameter $\delta\in(0,1)$, with probability at least $1-\delta$, we can obtain
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\bigl{(}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h-1}^{t}]\bigr{)}\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right).$
In the pessimistic case, it holds that $\alpha_{i}\leq\overline{\alpha}_{t,i}$ for every constraint $i$ and episode $t$. Thus, the bound on term (b) can be obtained simply by summing the estimation errors across the constraints. For terms (c) and (d), we can derive the bounds using the definitions and Lemma 21.

### v3

@@ 3.2 Estimators of Thresholds
We ask how to use the minimum amount of data while ensuring good performance in the stationary environment. To control computational cost without sacrificing statistical guarantees, we estimate thresholds using only the most recent episodes through a Growing-Window (GW) approach and then estimates at a data-efficient representative pair for each step.
Let $\{W_{t}\}_{t=1}^{T}$ be a window-length sequence that increases with episode $t$. For any constraint $i\in[m]$, step $h\in[H]$ and state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the GW estimator at episode $t$ is
$\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a):=\frac{\sum_{l=\max\{1,t-W_{t}+1\}}^{t}\tilde{\alpha}_{i,h}^{l}(s,a)\,\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}}{\max\{1,N_{h}^{(W_{t}),t}(s,a)\}},$
where $N_{h}^{(W_{t}),t-1}(s,a):=\sum_{l=\max\{1,t-W_{t}+1\}}^{t-1}\mathbf{1}_{\{s_{h}^{l}=s,a_{h}^{l}=a\}}$ counts visits to state-action pair $(s,a)$ in the last $W_{t}$ episodes before episode $t$. As $t$ increases, the window size $W_{t}$ grows. We prove its asymptotic consistency as follows.
@@
$\left|\hat{\alpha}_{i,h}^{(W_{t}),t}(s,a)-\alpha_{i,h}\right|\leq\zeta_{h}^{(W_{t}),t}(s,a),$
where $\zeta_{h}^{(W_{t}),t}(s,a)=\min\left\{1,\sqrt{\frac{4\ln(mSAHT/\delta)}{\max\{1,N_{h}^{(W_{t}),t-1}(s,a)\}}}\right\}$.
For brevity we simplify the notation $\hat{\alpha}_{i,h}^{(W_{t}),t}$ to $\hat{\alpha}_{i,h}^{t}$, $\zeta_{h}^{(W_{t}),t}$ to $\zeta_{h}^{t}$. The proof is given in Appendix C.
@@
Fixed-window approaches may repeatedly discard valuable historical data, which hurts long-run convergence. Full-history methods may respond too slowly to changes. A Growing-Window approach preserves enough past information to ensure eventual convergence but can adapt to evolving conditions more rapidly.
@@
Having defined the windowed counts and per-pair estimates, we now summarize each step $h$ with a statistically most informative representative pair drawn from the same window:
$(s_{h}^{\star},a_{h}^{\star})=\underset{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}{\arg\max}\;N_{h}^{(W_{t}),t}(s^{\prime},a^{\prime}).$
Using the representative rather than pooling all pairs avoids diluting the estimate with many low-frequency pairs, reduces computation and storage, and, within a stationary environment, retains asymptotic correctness as the window grows. We further consider two ‘extreme’ estimation strategies, namely pessimistic and optimistic versions for any $(s,a)$,
$\overline{\alpha}_{i,h}^{t}(s,a):=\hat{\alpha}_{i,h}^{t-1}(s,a)+\zeta_{h}^{t-1}(s,a),\quad\underline{\alpha}_{i,h}^{t}(s,a):=\hat{\alpha}_{i,h}^{t-1}(s,a)-\zeta_{h}^{t-1}(s,a),$
and evaluate them at the representative pair,
$\overline{\alpha}_{i,h}^{t}:=\overline{\alpha}_{i,h}^{t}(s_{h}^{\star},a_{h}^{\star}),\quad\underline{\alpha}_{i,h}^{t}:=\underline{\alpha}_{i,h}^{t}(s_{h}^{\star},a_{h}^{\star}).$ (2)
By construction, with probability $1-\delta$, any possible estimator $\hat{\alpha}_{i}$ satisfies inequality: $\underline{\alpha}_{i}\leq\hat{\alpha}_{i}\leq\overline{\alpha}_{i}$. These optimistic and pessimistic variants are the upper and lower confidence bounds for the GW estimator while maintaining the minimal-data design.
@@ 3.3 Feasibility and Strong Duality
A policy $\pi$ is defined to be pessimistic if the expected cumulative constraint is no less than the pessimistic threshold $\overline{\alpha}$; i.e., $V_{g}^{\pi}\geq\overline{\alpha}$. Let $\mathcal{F}_{\textup{pes}}$ denote the feasibility set,
@@
Similarly, we define a policy $\pi$ to be optimistic if the expected cumulative constraint is greater than or equal to the optimistic threshold $\underline{\alpha}$, i.e., $V_{g}^{\pi}\geq\underline{\alpha}$, and its feasibility set as below,
@@
From the definitions, pessimistic and optimistic policies are boundaries of all possible policies: $\mathcal{F}_{\textup{pes}}\subseteq\mathcal{F}_{q}\subseteq\mathcal{F}_{\textup{opt}}$ with probability $1-\delta$.
Throughout, we make the following assumptions, which is standard in the context of CMDPs altman1999constrained; efroni2020exploration; liu2021learning; liu2024learning.
Assumption 1 (Slater condition).
There exists $\pi^{0}\in\Pi$ and $\xi\in\mathbb{R}_{>0}^{m}$ such that $V_{g_{i}}^{\pi^{0}}\geq\alpha_{i}+\xi_{i}$ for all $i\in[m]$. Set the Slater gap $\rho=(r^{\top}q^{\pi^{\star}}(p)-r^{\top}q^{\pi^{0}})/\min_{i\in[m]}\xi_{i}$.
@@
Building on this assumption, we further require that the pessimistic tightening of the thresholds does not eliminate the existence of such a strictly feasible point. Similar ‘conservative tightening’ assumptions have appeared in recent work bura2022dope.
Assumption 2.
For $i\in[m]$, we have $\overline{\alpha}_{i}-\alpha_{i}<\xi_{i}$.
This condition ensures that the strictly feasible point from Assumption 1 remains feasible under the pessimistic setting. Their strong dualities are characterised by the following lemmas.
@@
Under Assumptions 1 and 2, we have $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{pes}}}\mathcal{L}(q,\lambda)=\max_{q\in\mathcal{F}_{\textup{pes}}}\min_{\lambda}\mathcal{L}(q,\lambda).$
@@
Under Assumption 1, we obtain $\min_{\lambda}\max_{q\in\mathcal{F}_{\textup{opt}}}\mathcal{L}(q,\lambda)=\max_{q\in\mathcal{F}_{\textup{opt}}}\min_{\lambda}\mathcal{L}(q,\lambda).$
Each lemma requires the existence of a policy that strictly satisfies the corresponding thresholds. We illustrate these feasible sets under pessimistic and optimistic thresholds and their corresponding feasible solutions in Figure 1. If there exists a policy that strictly satisfies pessimistic thresholds, for all possible thresholds, the strong duality holds.
@@ 3.4 A Primal-Dual Approach for CMDPs with Stochastic Thresholds
In this section, we propose SPOT (Stochastic Pessimistic-Optimistic Thresholding), a novel algorithm specifically tailored for CMDPs under stochastic thresholds. Unlike the closest algorithm OptPrimalDual-CMDP efroni2020exploration, which assumes a fixed and known threshold, SPOT tackles the more challenging setting in which the threshold is unknown, and potentially changing over time.
@@
0: episodes $T$, $\bf\lambda_{1}=0$, stepsizes $\eta_{t}$ and $\eta_{\lambda}$, and initial policy $\pi_{h}^{1}(a|s)=1/A$ for all $s$, $a$, $h$.
@@
9:  Execute $\pi_{t}$ via Algorithm 1 and update counters and empirical model (i.e., $N^{t},\hat{r}^{t},\hat{g}^{t},\hat{p}^{t},\hat{\alpha}^{t}$) via Equation (6)
@@
For a given dual variable, the algorithm first updates the primal policy and subsequently refines the dual value using the estimated value. In the first stage (Line 2), it updates the optimistic estimators $\overline{r}^{t}$ and $\overline{g}^{t}$. Then, it computes the estimator $\alpha_{\diamond}^{t}$ as prescribed in Equation (2). Importantly, depending on the problem context, estimator $\alpha_{\diamond}^{t}$ can be either pessimistic estimator $\overline{\alpha}^{t}$ or optimistic estimator $\underline{\alpha}^{t}$, thereby accommodating different degrees of uncertainty in the thresholds. Thus, at each episode $t$, it yields a new CMDP $\mathcal{M}_{t}:=(\mathcal{S},\mathcal{A},\overline{r}^{t},\overline{g}^{t},\alpha_{\diamond}^{t},\overline{p}^{t})$, which the algorithm then optimises via
$\max_{q}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}^{\prime}(q,\lambda):=\overline{r}^{\top}q+\lambda^{\top}(\overline{G}^{\top}q-\alpha_{\diamond}).$
@@
$\lambda_{t+1}=\Pi_{\mathcal{C}}\left(\lambda_{t}-\frac{1}{\eta_{\lambda}}(\overline{G}_{t-1}^{\top}q^{\pi_{t}}-\alpha_{\diamond}^{t})\right),$ (3)
where $\overline{G}_{t-1}^{\top}q^{\pi_{t}}$ is equal to $\hat{V}_{\overline{g}}^{\pi_{t}}$ and the projection set is defined as $\mathcal{C}:=\{\lambda\in\mathbb{R}^{m}:0\leq\lambda\leq\rho\bf 1\}$ with $\rho$ being the Slater gap defined in Section 3.3 and the projection operator $\Pi_{\mathcal{C}}:\mathbb{R}^{m}\to\mathcal{C}$ is defined by $\Pi_{\mathcal{C}}(x):=\arg\min_{y\in\mathcal{C}}\|y-x\|_{2}$. The feasibility proof for Equation (3) is presented in Appendix F. Finally, the policy $\pi_{t}$ is executed to collect new data, which is then used to update the empirical models and threshold estimates for subsequent iterations (Line 9).
@@
  • Stochastic and unknown thresholds vs. fixed thresholds. SPOT can handle thresholds that are unknown and random, going beyond the scope of fixed-threshold methods, which fit in more realistic, critical scenarios.
  • Pessimistic or optimistic thresholds balancing exploration and safety. SPOT allows using either pessimistic or optimistic thresholds, selected according to the application’s risk tolerance: pessimistic thresholds ensure robust feasibility under worst-case assumptions, while optimistic thresholds enable more ambitious exploration when additional risk is acceptable. This twofold approach empowers SPOT to adapt to varying levels of constraint uncertainty.
@@ 4.1 Reward Regret and Constraint Violation of a Pessimistic Policy
Theorem 2 (Bounds for reward regret and constraint violation of pessimistic policy).
Let $S$ be the number of states, $A$ be the number of actions, $H$ be the horizon length, $m$ be the number of constraints, $\rho$ be the Slater gap under Assumptions 1 and 2, $\gamma$ be the Growing-Window parameter, $T$ be the number of episodes and $\mathcal{N}$ denote the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability $1-\delta$, Algorithm 2 under pessimistic thresholds obtains the following regret bounds:
$\mathcal{R}_{T}(r)$ $\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right),$
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
The closest works would be Efroni et al. efroni2020exploration, who first introduce online methods that achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret and constraint violation in CMDPs; later, Ghosh et al. ghosh2024towards prove that $\tilde{\mathcal{O}}(\sqrt{T})$ regret and hard violation can be achieved by a primal-dual approach in tabular CMDPs. This paper removes the fixed-threshold assumption, and achieves comparable guarantees.
Remark 5.
Existing studies have also employed pessimistic thresholds, but the thresholds are predetermined bura2022dope. In contrast, our approach adaptively updates the thresholds during learning.
@@
We present a proof sketch here, with a detailed proof deferred to Appendix H.
@@
Decompose the sum of regret and violation scaled by dual variable, applying Lemmas 19 and 21, as follows,
$\mathcal{L}(\lambda)\leq\sum_{t=1}^{T^{\prime}}\left(f_{t}-\tilde{f}_{t}\right)+\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2},$ (4)
where $\mathcal{L}(\lambda)$ denotes the Lagrangian for any $0\leq\lambda\leq\rho\bf 1$ and $\eta_{\lambda}$ denotes the learning rate for dual update. The first term on the right side measures the gap between actual reward and the optimistic estimates; the second term quantifies the cumulative error between the estimated policy and the optimal policy with $\lambda_{t}$; the last two terms capture the error from updating the dual variables.
@@
Upper bound each term in the decomposition. For the first term and third term, we have $\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}\leq\frac{mH^{2}T}{2\eta_{\lambda}}$. By applying Lemma 20, we have $\left|\sum_{t=1}^{T^{\prime}}\left(f_{t}-\tilde{f}_{t}\right)\right|\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}\right)$. Bounding the second term is one of major technical contributions in this paper. We obtain the following lemma which is proved in Appendix G.
Lemma 3 (Policy optimality of pessimistic policy).
Conditioning on the success event (formally defined in Appendix D), for any episode $t\in[T^{\prime}]$, parameter $\gamma\in(0,1]$. Under pessimistic thresholds, we have
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left(\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}\right).$
@@
Let $\{\mathcal{F}_{h}^{t}\}_{(t,h)\in[T]\times[H]}$ be the filtration generated by the state-action sequences and threshold values observed up to time step $h$ at episode $t$. Suppose that for every $(t,h)$, $\zeta_{h}^{t}(s_{h},a_{h})$ satisfies $\left|\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}[\zeta_{h}^{t}(s_{h},a_{h})]|\mathcal{F}_{h}^{t-1}\right|\leq 2.$ Then, with probability $1-\delta$,
$\left|\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})-\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})|\mathcal{F}_{h}^{t-1}\right]\right|\leq\tilde{\mathcal{O}}\left(\sqrt{TH\text{ln}\frac{1}{\delta}}\right).$